OpenStudy (anonymous):

how do I solve : log(x+96) - log (2x+17) = log (x) ?

OpenStudy (anonymous):

use log(a)-log(b)=log(a/b) and log(c)=log(d) only when c=d

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problem, for your next question do similar, divide, and remember that 1=log_10 10

OpenStudy (anonymous):

thanks! I keep forgetting that!